OpenStudy (anonymous):

How can linear equations and functions be written and graphed? Explain how shifting the graph changes the function. MEDAL TO ANYBODY WHO HELP ME

OpenStudy (rhystic):

Linear equations are written as y = C1x + C2 where C1 & C2 are constants. They are graphed by plugging in numbers for y or x and solve for the other. For example if you have an equation y = 3x + 6 you then for y = 1 you have 1 = 3x + 6 -> x = -5/3 you then take the two coordinates (-5/3,1) and plot the point do this for further numbers and connect the points with a straight line. Shifting the graph up or down will change the value of C2 and shifting the graph left or right will change C1.

OpenStudy (anonymous):

OMG @Rhystic THANK YOU SO MUCH :)

OpenStudy (anonymous):

@Rhystic what are the key features of linear functions explain to find each function...